### v3

# What is Fair? Defining Fairness in Machine Learning for Health

Abstract
Ensuring that machine learning (ML) models are safe, effective, and equitable across all patient groups is essential for clinical decision-making and for preventing the reinforcement of existing health disparities. This review examines notions of fairness used in ML for health, including a review of why ML models can be unfair and how fairness has been quantified in a wide range of real-world examples. We provide an overview of commonly used fairness metrics and supplement our discussion with a case-study of an openly available electronic health record (EHR) dataset. We also discuss the outlook for future research, highlighting current challenges and opportunities in defining fairness in health.
Main
There are myriad potential applications of machine learning (ML) for health, including automated disease detection, computer-aided diagnosis, and personalized treatment planning [1]. However, there is substantial evidence that, without appropriate forethought and planning, ML models can introduce or exacerbate health inequities by making less accurate decisions for certain groups or individuals [2]. Within medical imaging, state-of-the-art ML models used for disease diagnosis, risk prediction, and triage management are known to underperform within minority groups defined by protected attributes, including sex, race, and ethnicity [3, 6, 7, 8, 9, 10, 11, 12, 13, 4, 5]. For example, deep learning models used to detect 14 common diseases from chest X-rays substantially were found to under-diagnose intersectional under-served subgroups, such as Hispanic female patients, potentially resulting in treatment delays if deployed in practice [11]. Similarly, ML models trained on electronic health record (EHR) and claims data have demonstrated biased performance by underperforming in specific subgroups [14, 15]. For example, in a landmark study, a widely used commercial risk prediction model for identifying patients with complex health needs was found to have significantly lower accuracy for Black patients compared to White patients, making Black patients less likely to be recommended for essential care services than equally healthy white patients [16].
As applications of ML in health become commonplace, it is crucial to recognize, account for, and mitigate such disparities in model performance in order to ensure health equity. Broadly, an ML model is said to be fair if it does not discriminate against an individual or group [17]. Concepts of fairness have been extensively studied across various disciplines, including social choice theory, game theory, economics, and law [18, 19, 20, 21, 22, 23, 24]. Building on these principles, the subfield of fairness within ML provides a framework for evaluating and mitigating bias throughout the model development process.
Although fairness has been an extremely active area of research over the past decade, evaluating and implementing fairness within ML for health is relatively nascent [25, 26, 27]. A systematic review of articles utilizing ML for EHR-based phenotyping found that only 5% of studies assessed fairness [28]. Similarly, a review of EHR-based prediction models found that most studies investigate the overall performance of ML models, but do not interrogate potential biases [29]. Beyond EHR applications, several scoping reviews of clinical ML models found that the adoption of fairness remains inconsistent, partly due to knowledge gap between ML and clinical researchers [27, 30, 31]. Moreover, over two dozen definitions of fairness have been proposed, most of which originate outside health-focused literature, making it a particularly challenging domain to navigate.
To review this fundamental topic, we examine key notions of fairness and their use in ML in health [^1] This review focuses on quantitative measures of fairness. Alternative approaches are discussed in the outlook section.. We first introduce biases that can emerge throughout the model development process using examples from the literature to illustrate why ML models can be unfair. We then review what it means for a model to be fair, beginning with an overview of the most common fairness framework, group fairness, and moving to the emerging frameworks of individual and causal fairness [32]. Our discussion includes the mathematical formulation of various fairness criteria, real-world examples, and a case study on mortality prediction using an openly available EHR dataset. We conclude with challenges in existing approaches and prospects for future research. Our review highlights that, in light of the large number of fairness notions, how to operationalize fairness in clinical settings remains an open problem. We therefore aim to provide a concise overview and organization of fairness definitions used within ML in health that can be used as a foundational resource for researchers navigating and contributing to the rich and nuanced fairness literature (see Supplementary Table S1 for a detailed comparison with existing references).
Why ML models are unfair
Even with a well-defined and well-intentioned research question, ML models can be unfair due to biases in the data, in the model, and/or in the deployment of the model [33, 34, 35, 36, 25, 37, 38, 39]. Figure 1 provides a visualization of sources of bias that can arise throughout the model development process and Table 1 details common sources of bias together with examples from the health-focused literature.
Figure 1: Sources of bias in developing a machine learning (ML) model. Bias can arise at each stage of the model development process, from data collection to deployment. Each stage influences the next, with bias being potentially perpetuated and compounded throughout the development process. After deployment, an unfair ML can also introduce or reinforce societal bias. This figure is adapted from [40] and [41].
Briefly expanding on Table 1, bias in the data arises from (i) societal and/or (ii) statistical bias [40]. With respect to the former, the data used in health applications most often measures and categorizes people and therefore encodes societal structures, injustices, and stereotypes, such as gender, racial, and age bias [42, 43]. For example, underlying social inequalities in healthcare access can limit the amount of data available in EHRs of certain subgroups, such as Black patients [44]. In terms of statistical bias, the data can also fail to represent the population of interest due the sampling method (selection bias), time of collection (temporal bias), or data quality issues such as missing, mismeasured, or insufficient data (missing data, measurement, and minority bias). In the fairness literature, data with any of these undesirable properties are informally referred to as biased [45, 17]. Observational clinical data (e.g., EHR, medical images) inherently contains various biases and researchers must carefully consider sources of bias in any application [46, 47]. Additionally, choices made during model training and evaluation can further amplify biases in the data or incorporate new bias, including selecting an inappropriate label (label bias), model (algorithmic bias), and/or evaluation metric or data set (evaluation bias). Table 1 details several health examples subject to model bias. Lastly, during deployment, bias in the data and/or model can be reinforced or introduced when users selectively disregard (dismissal bias) or overly trust (automation bias) a model’s outputs, or if models are inaccessible to certain subpopulation (privilege bias). For instance, ML-assisted diagnostic imaging requires that hospitals have the necessary equipment, yet fewer than 20% of rural emergency departments in Canada have in-house computed tomography (CT) scanners [48]. A review of biases in non-health applications can be found in [43] and [34].
Dr. Leo Celi, a health AI expert, emphasizes that “data bias is the waterloo of health AI” and argues that bias must be a primary consideration throughout the ML pipeline - from problem formulation to model deployment [49]. Building on existing AI reporting guidelines, [25] introduced a bias evaluation checklist to enable practitioners to systematically and holistically address bias in clinical predictive models [50, 51]. Bias mitigation strategies and common pitfalls in applying ML in health are further discussed in [52] and [53]. The fairness criteria we introduce in the subsequent sections play an important role in this process. The criteria are typically integrated directly into model training or used as evaluative metrics to identify disparities after the model has been trained [54, 55, 56, 57, 41, 58]. While this review focuses on how to define fairness, strategies for mitigating bias are an equally active and important area of research and have been reviewed from a methodological perspective in [59].
(a) Bias in the data
Type of Bias Definition Example
Societal/historical Bias The data reflects long-standing societal disparities encoded within the data over time. Clinical word embeddings trained on large corpora of text, such as clinical notes from healthcare systems, reflect biases about ethnic minorities [60].
Selection Bias The data is not representative of the population of interest. Data in volunteer-based biobanks does not reflect the general population as participation rates are higher among women, older age groups, and individuals with higher socioeconomic status. [61].
Measurement Bias The data contains variables that are collected or measured inaccurately. Gender identity is often recorded in EHRs without provider education and therefore may not accurately represent an individual’s identity [62].
Temporal Bias The data captures a specific time period that may not reflect current or future conditions. Pediatric mental health visits were lower than expected for individuals with lower socioeconomic status during the first year of the COVID-19 pandemic [63].
Minority Bias The data lacks adequate representation from the minority group for the model to accurately learn. Most of the data for genetic studies is from European ancestry populations [64].
Missing Data Bias The data has variables that are incompletely measured. Patients from low-income backgrounds have higher rates of missing medical measurements, which is often due to inequities in access to healthcare [65].
(b) Bias in the model
Type of Bias Definition Example
Label Bias An imperfect proxy is selected to train a model instead of the outcome of interest. Healthcare cost was used as a proxy for healthcare need in a commercial algorithm used to identify patients for high-risk care management programs. The algorithm significantly under-identified Black patients for care services as less money is spent on Black patients relative to similarly healthy white patients [16].
Algorithmic Bias Properties of a model and/or its training algorithm create or amplify bias in the data. Differentially private prediction models for mortality performed poorly for minority groups [66].
Evaluation Bias An inappropriate choice of benchmark data or metrics are used for evaluation. Existing public skin disease AI benchmarks do not have images of biopsy-proven malignancy, the gold standard for disease annotation, on dark skin [67, 68].
(c) Bias in deployment
Type of Bias Definition Example
Automation Bias Model users overly trust model outputs, sometimes even against their own knowledge. A computer-aided diagnosis system that scans mammograms and marks suspicious areas of potential cancer features had lower sensitivity for women aged 40–49 compared to older age groups. A clinician using this system can disproportionately miss cancers for this age group [69].
Dismissal Bias Model users ignore model recommendations, often due to frequent false alerts. If the same computer-aided diagnosis system also has a lower positive predictive value for women aged 40–49, then clinicians may disregard its recommendations, as they are more likely to be false positives [70].
Privilege Bias Models are not available or cannot be deployed in all settings, such as parts of the world where medical technology or sensors are unavailable. Less than 20% of rural emergency departments in Canada have in-house access to computed tomography (CT) scanners [48].
Table 1: Common biases that arise in machine learning (ML) applications in health and an example from the literature.
What it means for a ML model to be fair
Existing definitions of fairness primarily fall into three categories: group fairness, individual fairness, and causal fairness. Group fairness criteria are commonly used in health and deem a model as fair if its predictions are similarly accurate or calibrated across a predefined set of groups. These groups are most often defined by a protected attribute(s) such as age or race [^2] Sensitive attribute and protected attribute are used interchangeably [71]. Fairness is also evaluated across groups defined by social determinants of health (e.g., income and education) [72]. [34]. Common protected attributes used in the health-focused literature are summarized in Table 2. In contrast, individual fairness is a less commonly used framework and requires that the model provide similar predictions to similar individuals based on a user-defined similarity metric [24, 31]. Lastly, causal fairness criteria utilize causal estimands to quantify unfairness and link observed disparities in model performance to their underlying cause [73, 74]. Causality-based fairness definitions are particularly attractive for health-focused applications as they enable practitioners to interrogate biases [75].
Attribute Examples
Age [9, 76, 77, 78]
Disability [76]
Marital Status [79]
National Origin [78]
Race [9, 79, 76, 77, 80, 78, 33, 25]
Sex [9, 79, 76, 80, 78]
Socioeconomic Status [9, 79, 77, 78, 33, 25]
Table 2: Commonly used protected attributes in machine learning for health. Examples of studies that evaluate fairness with respect to each protected attribute.
Group fairness criteria are referred to as “oblivious” as they equate fairness with parities based solely on the distribution of the data (i.e., the predictions, outcome, and protected attribute), while individual and causal fairness criteria are “nonoblivious” as they require additional context in the form of user-defined similarity metrics and causal models, respectively. Importantly, many fairness criteria are incompatible in that they cannot be simultaneously satisfied. A taxonomy of fairness, including different notions of group, individual, and causal fairness and their incompatibilities that are introduced in subsequent sections, is presented in Figure 2.
Figure 2: Taxonomy of fairness. Fairness criteria primarily fall into three categories: group, individual, and causal fairness. Group fairness criteria are oblivious in the sense that they can be entirely inferred from the distribution of the data. Individual and causal fairness criteria are non-oblivious criteria as they require specification of a similarity metric and a causal model, respectively. Many fairness criteria cannot be simultaneously satisfied and incompatibilities between different notions of fairness are depicted with a red X. This taxonomy is not exhaustive and includes the notions of fairness introduced in this review.
Notation
Throughout, we denote the outcome of interest as $Y$, the features used for model training as ${\bf X}$ [^3] Note that ${\bf X}$ does not have to be a vector; it could be an image, text, tensor, or any other type of data., and the categorical variable for the protected attribute as $A$. ${\bf X}$ may or may not contain $A$ and we discuss this issue when we introduce individual fairness. We let $S=\hat{f}({\bf X})$ be the output from the ML model, where $\hat{f}$ is learned from a set of training data. For example, in classification, $Y$ is a binary label for membership in the positive or negative class and $S$ is the predicted probability of being in the positive class, referred to as the score. The final decision, $D$, is positive if $S$ exceeds a pre-determined classification threshold and negative otherwise.
Group fairness
Group fairness criteria require ML models to perform similarly across groups of interest and are the most popular fairness framework in health-focused applications [32]. The criteria primarily fall into three categories: independence, separation, and sufficiency [81, 17]. Table 3 provides mathematical definitions of various criteria and their interpretation in the context of an example. Our discussion focuses on classification as it is most widely used and well-studied [40, 34, 82]. Approaches for continuous outcomes, continuous protected attributes, categorical protected attributes that define more than two groups, and multiple protected attributes (i.e., subgroup fairness) are discussed in Supplementary Sections 2.
Independence
Under independence, an ML model is said to be fair if its decisions do not depend on the protected attribute (i.e., $D\perp\!\!\!\perp A$). Statistical (or demographic) parity is a common measure of independence that requires that the model classify individuals into the positive class at the same rate in each group [24, 83]. Conditional statistical parity relaxes this concept by requiring the rate of positive classifications to be the same within more granular groups defined by the protected attribute and other relevant factors. Independence-based metrics are infrequently used in health-focused applications as the prevalence of clinical outcomes often differs across groups (e.g., heart disease in males and females). Enforcing independence may prevent a model from learning a genuine association between the protected attribute and the outcome, which can decrease the overall performance of the model for the entire population [84, 85].
Separation
Separation requires that the model’s decisions do not depend on the protected attribute within the positive and negative classes (i.e., $D\perp\!\!\!\perp A\mid Y)$. This implies that, among individuals in the positive (or negative) class, the rate of making a positive (or negative) decision is consistent across groups. Common separation-based metrics therefore aim to equalize error rates across the groups, including the false negative rate (FNR, known as equal opportunity), false positive rate (FPR, known as predictive equality), or both (known as equalized odds).
These metrics have been widely used in health-focused applications as incorrect decisions can have severe consequences. The choice of a separation-based metric, however, depends on the specific clinical context. When false negatives have the most severe consequences, equal opportunity may be preferred. For instance, equal opportunity was used in a study of state-of-the-art computer vision models used to detect common diseases from chest X-rays, where a false negative corresponded to incorrectly identifying a patient as not having “no finding” on their X-ray [11]. The models had higher FNRs in several intersectional under-served subpopulations, such as Hispanic female patients, potentially resulting in delayed access to care. There are also situations where balancing both FNR and FPR is more appropriate and equalized odds should be prioritized [54]. For instance, in the context of rapid screening tools for COVID-19 used in emergency departments, a high FNR can lead to insufficient monitoring of COVID-19 while a high FPR can result in unnecessary additional testing [86]. Additional separation-based metrics are detailed in Table 3, including balance for the positive class and balance for the negative class [87].
Sufficiency
Sufficiency aims to equalize error rates among individuals with similar decisions [82]. Formally, sufficiency requires that the label does not depend on the protected attribute given the model’s decision ( i.e., $Y\perp\!\!\!\perp A\mid D)$. The decision is therefore “sufficient” for predicting the outcome in the sense that it subsumes the protected attributes [17]. Common sufficiency-based metrics focus on equalizing the positive predictive value (PPV, known as predictive parity), both the PPV and negative predictive value (NPV, known as conditional use accuracy equality), and calibration (known as well-calibration )[45, 81, 87, 88].
Sufficiency-based metrics are commonly used to evaluate models for disease screening, management, and triage. Conditional use accuracy equality is a particularly common metric, as disparities in PPV and NPV can have distinct impacts on patient care [89, 86]. For example, in COVID-19 triage algorithms, conditional use accuracy seeks to ensure that ICU allocations based on the model’s decisions are determined by medical needs rather than a protected attribute. Balancing the model’s PPV across groups means that patients prioritized for ICU care due to potential severe complications (e.g., high-risk cases) have an equal likelihood of actually experiencing these complications, regardless of protected attributes. Similarly, balancing the NPV across groups ensures that low-risk cases have an equal likelihood of not experiencing complications. Calibration-based metrics are also frequently applied. Well-calibration was used in our introductory example of the commercial risk prediction algorithm used to enroll patients in a high-risk care management program. In this example, healthcare costs were used as a proxy for health needs in training the model. As a result, at equivalent model scores, Black patients were in significantly poorer health than white patients as historically less money had been spent on their healthcare, meaning that they had to be sicker to qualify for the program [16].
Metric Definition Interpretation Relax-
ation
Independence-based criteria
Statistical Parity $P(D=1\mid A=a_{0})=P(D=1\mid A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ In predicting hospitalization and emergency department (ED) visits in heart failure, the predicted probability of a hospitalization or an ED visit is the same for males and females [90].
Conditional Statistical Parity $P(D=1\mid{\bf Z},A=a_{0})=P(D=1\mid{\bf Z},A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ In predicting hospitalization and emergency department (ED) visits in heart failure, the predicted probability of a hospitalization or an ED visit is the same for males and females after adjusting for age and pre-existing conditions [90].
Separation-based criteria
Equalized Odds $P(D=1\mid Y=0,A=a_{0})=P(D=1\mid Y=0,A=a_{1})\quad\text{and}$
$P(D=0\mid Y=1,A=a_{0})=P(D=0\mid Y=1,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$
In predicting asthma exacerbation in children, the rates of false positives (children without exacerbation incorrectly identified as having exacerbation) and false negatives (children with exacerbation incorrectly identified as not having exacerbation) are the same for children from low and high socio-economic classes [91].
Predictive Equality $P(D=1\mid Y=0,A=a_{0})=P(D=1\mid Y=0,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ A model predicting cardiovascular disease (CVD) diagnoses CVD in males and females without CVD at the same rate [92].
Equal Opportunity $P(D=0\mid Y=1,A=a_{0})=P(D=0\mid Y=1,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ A model predicting a suspicious finding from a chest X-rays misses findings in males and females at the same rate [93].
Balance for Positive Class $\mathbb{E}(S\mid Y=1,A=a_{0})=\mathbb{E}(S\mid Y=1,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ In a model used to screen for lung cancer from chest X-rays, the average predicted score of people with lung cancer is equal for males and females [93].
Balance for Negative Class $\mathbb{E}(S\mid Y=0,A=a_{0})=\mathbb{E}(S\mid Y=0,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ In a model used to screen for lung cancer from chest X-rays, the average scores among males and females without lung cancer are equal [93].
Sufficiency-based criteria
Conditional Use Accuracy Equality $P(Y=1\mid D=1,A=a_{0})=P(Y=1\mid D=1,A=a_{1})$ and
$P(Y=0\mid D=0,A=a_{0})=P(Y=0\mid D=0,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$
In predicting hospital readmission, the probability of being readmitted given the model makes that prediction and the probability of not being readmitted given the model makes that prediction is the same for Black and white patients [94].
Predictive Parity $P(Y=1\mid D=1,A=a_{0})=P(Y=1\mid D=1,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ In predicting hospital readmission, the probability of being readmitted given the model makes that prediction is the same for Black and white patients [94].
Well Calibration $P(Y=1\mid S=s,A=a)=s\quad\forall a\in\mathcal{A},s\in[0,1]$ In a model used to predict in-hospital mortality, the predicted event rates match the observed event rates at all values of the score for males and females [95].
Test Fairness $P(Y=1\mid S=s,A=a_{0})=P(Y=1\mid S=s,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A},s\in[0,1]$ In a model used to predict in-hospital mortality, the predicted event rates are the same for males and females at all values of the score [95].
Other criteria
Brier Score Parity $\mathbb{E}[(Y-S)^{2}\mid A=a_{0}]=\mathbb{E}[(Y-S)^{2}\mid A=a_{1}],\quad\forall a_{0},a_{1}\in\mathcal{A}$ In a model used to enroll patients into high-risk care management programs, the mean squared error between the score and the label is the same for Black and White patients [16].
Overall Accuracy Equality $P(D=Y\mid A=a_{0})=P(D=Y\mid A=a_{1})\quad\forall a_{0},a_{1}\in{\cal A}$ In a model used to screen for lung cancer from chest X-rays, the probability of a correct classification (i.e., correctly classifying a patient as having or not having lung cancer) is the same for males and females [93].
Treatment Equality #$\frac{\#\{D=0,Y=1,A=a_{0}\}}{\#\{D=1,Y=0,A=a_{0}\}}=\frac{\#\{D=0,Y=1,A=a_{1}\}}{\#\{D=1,Y=0,A=a_{1}\}}\quad\forall a_{0},a_{1}\in{\cal A}$ In a model used to screen for lung cancer from chest X-rays, the ratio of the false negatives and false positives is the same for both males and females [93].
Table 3: Common group fairness criteria. Mathematical definitions of group fairness criteria, their interpretation in the context of an example from the literature, and an indication of whether the criterion is a relaxation of independence ($D\perp\!\!\!\perp A$), separation ($D\perp\!\!\!\perp A\mid Y$), or sufficiency ($Y\perp\!\!\!\perp A\mid D$). Symbols: #=\#=# = number of, $P=$ probability, $E=$ expected value. Notations: $Y:$ outcome, ${\bf X}:$ features used for model training, $A:$ protected attribute that takes value in the set $\mathcal{A}$, $S:$ model score, $D:$ model classification based on thresholding $S$, ${\bf Z}$: additional set of features.
Incompatibilities
Independence, sufficiency, and separation provide different perspectives on what it means for a model to be fair. Except under highly restrictive conditions, it is not possible for an algorithm to fulfill all criteria simultaneously [54, 45, 96]. It is therefore critical for healthcare researchers to choose which group fairness considerations are most relevant to the specific context. More specifically, the following pairs of criteria are incompatible in the sense that they cannot generally be simultaneously satisfied: independence and sufficiency, independence and separation, and separation and sufficiency. A basic requirement for any of these pairs to hold is that the outcome and the protected attribute are marginally independent (i.e., $Y\perp\!\!\!\perp A$). In classification problems, this means the probability of being in the positive class is the same across groups - a condition often violated in clinical contexts, where disease prevalence may vary by factors like age or sex. Supplementary Section 3 provides additional mathematical details related to the incompatibilities and also introduces approximate fairness, which allows for small deviations in group fairness metrics to ease these incompatibilities. In Supplementary Section 4, we further explore the practical implications of these incompatibilities through a case study using an openly available EHR dataset.
Individual fairness
In contrast to group fairness that targets the average performance of a model across groups, individual fairness ensures that “like cases are treated alike”, an idea grounded in Aristotle’s consistency principles [24, 97]. One of the first works on individual fairness, fairness through awareness (FTA), formalized the notion of similarity by defining metrics that measure the distance between individuals and their predictions. Fairness is achieved if the distance between the prediction for any two individuals is at most the distance between them. A special case of FTA is when the model yields identical decisions for two individuals with all but their protected attributes being the same, a property known as causal discrimination [98].
The implementation of FTA-based approaches critically depends on the choice of distance metrics defined on both predicted outcomes and individuals. For predictions, the choice varies by the type of outcome under study. For example, the Euclidean distance is used for logit transformed predicted probabilities in classification while the absolute difference of the hazard function has been used for survival analysis [99, 100]. However, the metric used to measure the similarity between individuals has been a focus of ongoing research as it encodes exactly how individuals are similar, which relies on an “awareness” of the context [^4] For consistency index proposed by [101], the distance metric for predictions are absolute difference, but it still relies on the choice of metric to define $k$ nearest neighborhood.. In the original FTA paper, the authors suggest leaving the choice of metric to domain experts. As there is currently no agreed upon standard, particularly within health settings, various strategies have been proposed to learn the metric from available data [102, 103, 104]. A scoping review on individual fairness in healthcare provides a comprehensive list of individual fairness methods and corresponding software [31]. The review highlights that individual fairness is just emerging within health-focused applications, likely due to the relative nascency of individual fairness methods developed within the ML literature and the difficult in appropriately quantifying similarity among individuals.
It is important to note that FTA is a counter approach to fairness through unawareness (FTU), which is intended to be a catch all solution to prevent bias by not explicitly including protected attributes into modeling [105] [^5] Note that fairness through awareness implies causal discrimination, but not vice versa.. However, removing protected attributes from a model warrants careful consideration in health-focused applications. First, protected attributes may serve as critical predictors for the outcome of interest, such as age in sepsis or cardiovascular disease prediction and race in cancer screening models [106, 107, 108, 109]. Excluding these predictors can diminish the overall predictive accuracy and impact all individuals adversely or even lead to bias against the majority group. Second, protected attributes are often highly correlated with non-protected attributes. Simply removing protected attributes from the model does not prevent the model from inferring them from other attributes. This is evident in a recent study that found ML models can infer race from medical images [110]. Finally, the quality of the other predictors included in the model may vary across the protected groups (e.g., rates of missing data, measurement error). For instance, in cancer risk prediction, family history serves as an important predictor for disease risk, but was shown to be less reliably documented for Black participants in EHR and therefore less predictive for models developed with these data. [111] found that including race as a predictor significantly improved model performance compared to race-blind algorithms for colorectal cancer. FTU is therefore not always ethical or achievable [112].
Lastly, individual fairness is motivated by the inherent weakness in group fairness criteria that only consider average model performance within groups. That is, there are situations in which group fairness can be satisfied, but individuals within a group can be discriminated against. The incompatibilities among individual and group fairness have been examined from a theoretical perspective in [113, 114, 82, 115]. For example, in predicting ICU admission, if the rate of admission differs significantly across protected attributes, the predictions must be adjusted to achieve statistical parity and obtain an equal rate of positive predictions across the groups. This can cause individuals with similar clinical profiles and medical needs to receive different predictions based solely on their protected attributes, which contradicts the principle of individual fairness. However, individual fairness implies statistical parity when the Wasserstein distance between two groups’ protected attribute distributions is small [24, 116].
Causal fairness
We conclude with a brief overview of causal fairness, an increasingly prominent topic in ML that warrants a dedicated review within the health context [117, 118, 119, 120, 121, 31, 75]. As the name suggests, causal fairness focuses on understanding the causal relationships between protected attributes and a model’s decisions [30]. Although causality-based fairness approaches are especially valuable in health applications for mitigating confounding effects and disentangling mechanisms of bias, the use of causal fairness is health is in its early stages.
We introduce one way of incorporating causality into fairness through counterfactuals, which are simply “what-if” statements [122, 123]. In fairness context, the counterfactual is often the unobserved model’s decision that would have happened if the protected attribute had been different [105, 118, 124]. One of the first counterfactual-based fairness criteria, simply called (individual) counterfactual fairness, deems a model fair if the observed and counterfactual predictions align for an individual when their protected attributed is changed, given all other variables of interest are held constant [105] [^6] Counterfactual fairness is also viewed as an individual fairness criteria as it is grounded in achieving similar treatment for similar individuals[105].. To better explain this concept, consider an over-simplified view of the world in predicting 30-day hospital readmission. The directed acyclic graph (DAG) in Figure 3 illustrates that the model’s decision for 30-day hospital readmission is directly related to previous diagnoses, age, and prescriptions, while race indirectly influences the decision through its impact on access to healthcare, which in turn influences historical diagnoses [125]. When predicting 30-day hospital readmission, counterfactual fairness is satisfied if a Black patient receives the same prediction they would have received if they were counterfactually of a different race, given all other variables are held constant. The ceteris paribus condition here is crucial, as in reality healthcare utilization may differ across racial communities, which could influence the rate of 30-day hospital readmission in practice. A challenge to causal fairness criteria is that they require a working causal model, the assumptions of which cannot be verified from observational data [105, 126, 118].
Other quantities from the causal inference literature have since been proposed, such as counterfactual parity which measures a population-level, rather than a individual-level, causal effect of the protected attribute on the model’s decision [127, 40, 128]. Conditional counterfactual fairness is a weaker criteria that considers the total effect of the protected attribute on the decision conditional on a set of variables of interest [105]. In our previous example, this means that for each group of individuals with identical characteristics (e.g., those with the same historical diagnosis, age and prescriptions), the 30-day hospital readmission does not change if their race were counterfactually altered. Another criteria used in health-focused applications is principle fairness, which is a population-level fairness criteria that evaluates decisions among individuals who would be similarly impacted by a model’s decision [129]. Formally, a model satisfies principle fairness if the model’s decision is independent of the protected attribute in principle stratum, which are subgroups with the same values of the counterfactual decision under all possible decisions [129]. More concretely, in our example, principle fairness is achieved if Black and white people have an equal chance of being flagged for 30-day readmission if they have the same likelihood of 30-day readmission (i.e., they belong to the same principal strata). These definitions, along with extensions of the previously introduced group fairness measures — such as counterfactual equalized odds, counterfactual predictive parity, and conditional principal fairness — are formally introduced in Supplementary Section 6. Further work on causal fairness, such as path-specific fairness, direct and indirect discrimination, and interventional fairness, are detailed in [130, 131, 132, 133, 134, 135, 136, 126, 137].
Figure 3: Directed acyclic graph (DAG) for predicting 30-day hospital readmission. The DAG illustrates the relationships among various factors influencing the predicted 30-day readmission. Historical diagnoses, age, and prescriptions have a direct influence on predicted 30-day readmission. Race indirectly affects the prediction through access to care, which in turn influences historical diagnoses.
Outlook
This review explores why models are unfair and various ways fairness has been defined in ML for health. Despite substantial progress in recognizing the importance of bias and in applying group, individual, and causal-based criteria in health applications, there is no consensus on how to appropriately quantify fairness [138, 139]. Each framework comes with unique challenges. Group fairness criteria are easy to implement and interpret, but limited by potential individual fairness violations and by incompatibilities that exist among various criteria. Likewise, individual fairness criteria prioritize similar outcomes for similar individuals to enhance equity in decision making, but can mask disparities at the group level and leave certain populations underserved. Causal fairness, which is gaining traction in health, enables researchers to investigate root causes of unfairness to inform fair decision-making. However, specifying reliable causal models can be challenging in health contexts. Balancing these different frameworks involves weighing the benefits of group-level equity, individual needs, and causal insights [140]. Researchers must remain informed about the full range of available methods and work within interdisciplinary teams (e.g., data scientists, clinicians, ethicists) to make appropriate decisions for their specific application.
It is critical to note that a quantitative understanding of fairness is just one component of operationalizing fairness in ML for health. Establishing common principles, guidelines, and standards that emphasize fairness across the entire model development process is essential, as simply assessing fairness after model training is insufficient [141]. Moreover, most existing fairness definitions were developed outside the health context and do not account for underlying inequities or potential harms to underserved populations [47]. There is an urgent need for context-aware fairness criteria tailored to the unique challenges of health applications [142].
Additionally, health data present inherent challenges, making it difficult - or sometimes impossible - to measure fairness accurately [143, 144, 145, 146, 147, 148]. Various protected attributes (e.g., attributes such as religion, gender identity, and sexual orientation that are missing from Table 2) are often unavailable or inaccurately measured in health datasets due to socio-cultural challenges and lack of standardized data collection practices [149, 62, 150]. While by no means a remedy for inadequate data, there has been recent progress in fairness approaches that operate with limited data on protected attributes that may prove useful in health applications [151, 152]. To date, however, many protected attributes have been omitted from fairness considerations, which underscores the necessity of future methodological work to acknowledge “a multiplicity of considerations, from privacy preservation, context sensitivity and process fairness, to an awareness of sociotechnical impact and the increasingly important role of inclusive and participatory research processes” [153]. We anticipate future work in these aforementioned directions, particularly in strategies focused on data collection and promoting data equity in health contexts, to establish a strong foundation for fair decision-making.
Acknowledgements
J. Gronsbell is grateful for support of an NSERC Discovery Grant (RGPIN-2021-03734) and a University of Toronto Data Science Institute Seed Funding for Methodologists Grant.
Author contributions
JeG conceived and supervised the study. JeG and JiG drafted the manuscript. JiG and BC conducted the case study and developed the R package. PV provided clinical oversight, ZM and CH provided additional guidance on ML, and HT provided ethical oversight. All authors provided valuable feedback on the manuscript.

### v4

@@
-What is Fair? Defining Fairness in Machine Learning for Health
+references.bib
 Abstract
@@
 Main
-There are myriad potential applications of machine learning (ML) for health, including automated disease detection, computer-aided diagnosis, and personalized treatment planning [1].
-However, there is substantial evidence that, without appropriate forethought and planning, ML models can introduce or exacerbate health inequities by making less accurate decisions for certain groups or individuals [2].
-Within medical imaging, state-of-the-art ML models used for disease diagnosis, risk prediction, and triage management are known to underperform within minority groups defined by protected attributes, including sex, race, and ethnicity [3, 6, 7, 8, 9, 10, 11, 12, 13, 4, 5].
-For example, deep learning models used to detect 14 common diseases from chest X-rays substantially were found to under-diagnose intersectional under-served subgroups, such as Hispanic female patients, potentially resulting in treatment delays if deployed in practice [11].
-Similarly, ML models trained on electronic health record (EHR) and claims data have demonstrated biased performance by underperforming in specific subgroups [14, 15].
-For example, in a landmark study, a widely used commercial risk prediction model for identifying patients with complex health needs was found to have significantly lower accuracy for Black patients compared to White patients, making Black patients less likely to be recommended for essential care services than equally healthy white patients [16].
-As applications of ML in health become commonplace, it is crucial to recognize, account for, and mitigate such disparities in model performance in order to ensure health equity.
-Broadly, an ML model is said to be fair if it does not discriminate against an individual or group [17].
-Concepts of fairness have been extensively studied across various disciplines, including social choice theory, game theory, economics, and law [18, 19, 20, 21, 22, 23, 24].
+There are myriad potential applications of machine learning (ML) for health, including automated disease detection, computer-aided diagnosis, and personalized treatment planning [rajpurkar2022ai].
+However, there is substantial evidence that, without appropriate forethought and planning, ML models can introduce or exacerbate health inequities by making less accurate decisions for certain groups or individuals [grote_enabling_2022].
+Within medical imaging, state-of-the-art ML models used for disease diagnosis, risk prediction, and triage management are known to underperform within minority groups defined by protected attributes, including sex, race, and ethnicity [rajpurkar_chexnet_2017, ting_development_2017, liu_comparison_2019, lee_machine_2020, seyyed-kalantari_chexclusion_2020, larrazabal_gender_2020, seyyed-kalantari_underdiagnosis_2021, puyol-anton_fairness_2022, bressem_deep_2021, roest_comparative_2022, lin_evaluate_2023].
+For example, deep learning models used to detect 14 common diseases from chest X-rays substantially were found to under-diagnose intersectional under-served subgroups, such as Hispanic female patients, potentially resulting in treatment delays if deployed in practice [seyyed-kalantari_underdiagnosis_2021].
+Similarly, ML models trained on electronic health record (EHR) and administrative claims data have demonstrated biased performance by underperforming in specific subgroups [li_evaluating_2023, banda_characterizing_2023].
+In a landmark study, a widely used commercial risk prediction model for identifying patients with complex health needs had significantly lower accuracy for Black patients than for White patients.
+This disparity made Black patients less likely than White patients with the same level of illness to be recommended for essential care services [obermeyer_dissecting_2019].
+As applications of ML in health become commonplace, it is crucial to recognize, account for, and mitigate such disparities in model performance to ensure health equity.
+Broadly, an ML model is said to be fair if it does not discriminate against an individual or group [barocas_fairness_2023].
+Concepts of fairness have been extensively studied across various disciplines, including social choice theory, game theory, economics, and law [young_equity_1995, roemer_equality_1998, roemer_theories_1998, rawls_theory_1971, rawls_justice_2001, hutchinson201950, dwork_fairness_2012].
 Building on these principles, the subfield of fairness within ML provides a framework for evaluating and mitigating bias throughout the model development process.
-Although fairness has been an extremely active area of research over the past decade, evaluating and implementing fairness within ML for health is relatively nascent [25, 26, 27].
-A systematic review of articles utilizing ML for EHR-based phenotyping found that only 5% of studies assessed fairness [28].
-Similarly, a review of EHR-based prediction models found that most studies investigate the overall performance of ML models, but do not interrogate potential biases [29].
-Beyond EHR applications, several scoping reviews of clinical ML models found that the adoption of fairness remains inconsistent, partly due to knowledge gap between ML and clinical researchers [27, 30, 31].
+Although fairness has been an extremely active area of research over the past decade, implementing fairness within ML for health is relatively nascent [wang_bias_2022, feng_fair_2022, liu_translational_2023].
+A systematic review of articles utilizing ML for EHR-based phenotyping found that only 5% of studies assessed fairness [yang_machine_2023].
+Similarly, a review of EHR-based prediction models found that most studies investigate the overall performance of ML models, but do not interrogate potential biases [chen2024unmasking].
+Beyond EHR applications, several scoping reviews of clinical ML models found that the adoption of fairness remains inconsistent, partly due to a knowledge gap between ML and clinical researchers [liu_translational_2023, plecko_causal_2022, anderson2024algorithmic].
 Moreover, over two dozen definitions of fairness have been proposed, most of which originate outside health-focused literature, making it a particularly challenging domain to navigate.
@@
 We first introduce biases that can emerge throughout the model development process using examples from the literature to illustrate why ML models can be unfair.
-We then review what it means for a model to be fair, beginning with an overview of the most common fairness framework, group fairness, and moving to the emerging frameworks of individual and causal fairness [32].
+We then review what it means for a model to be fair, beginning with an overview of the most common fairness framework, group fairness, and moving to the emerging frameworks of individual and causal fairness [awasthi2020beyond].
 Our discussion includes the mathematical formulation of various fairness criteria, real-world examples, and a case study on mortality prediction using an openly available EHR dataset.
@@
 Why ML models are unfair
-Even with a well-defined and well-intentioned research question, ML models can be unfair due to biases in the data, in the model, and/or in the deployment of the model [33, 34, 35, 36, 25, 37, 38, 39].
-Figure 1 provides a visualization of sources of bias that can arise throughout the model development process and Table 1 details common sources of bias together with examples from the health-focused literature.
-Figure 1: Sources of bias in developing a machine learning (ML) model.
+Even with a well-defined and well-intentioned research question, ML models can be unfair due to biases in the data, in the model, and/or in the deployment of the model [rajkomar_ensuring_2018, mehrabi_survey_2021, fazelpour_algorithmic_2021, mhasawade2021machine, wang_bias_2022, xu_algorithmic_2022, drukker_toward_2023, ferrara_fairness_2024].
+Figure 1 provides a visualization of sources of bias that can arise throughout model development and Table 1 details common biases using examples from the literature.
+Figure 1: Sources of bias.
 Bias can arise at each stage of the model development process, from data collection to deployment.
 Each stage influences the next, with bias being potentially perpetuated and compounded throughout the development process.
-After deployment, an unfair ML can also introduce or reinforce societal bias.
-This figure is adapted from [40] and [41].
-Briefly expanding on Table 1, bias in the data arises from (i) societal and/or (ii) statistical bias [40].
-With respect to the former, the data used in health applications most often measures and categorizes people and therefore encodes societal structures, injustices, and stereotypes, such as gender, racial, and age bias [42, 43].
-For example, underlying social inequalities in healthcare access can limit the amount of data available in EHRs of certain subgroups, such as Black patients [44].
+After deployment, an unfair model can also introduce or reinforce societal bias.
+This figure is adapted from [mitchell_algorithmic_2021] and [mbakwe2023fairness].
+Briefly expanding on Table 1, bias in the data arises from (i) societal and/or (ii) statistical bias [mitchell_algorithmic_2021].
+With respect to the former, the data used in health applications most often measures and categorizes people and therefore encodes societal structures, injustices, and stereotypes, such as gender, racial, and age bias [iliadis2016critical, suresh2021framework].
+For example, underlying social inequalities in healthcare access can limit the amount of data available in EHRs of certain subgroups, such as Black patients [getzen_mining_2023].
 In terms of statistical bias, the data can also fail to represent the population of interest due the sampling method (selection bias), time of collection (temporal bias), or data quality issues such as missing, mismeasured, or insufficient data (missing data, measurement, and minority bias).
-In the fairness literature, data with any of these undesirable properties are informally referred to as biased [45, 17].
-Observational clinical data (e.g., EHR, medical images) inherently contains various biases and researchers must carefully consider sources of bias in any application [46, 47].
+In the fairness literature, data with any of these undesirable properties are informally referred to as biased [chouldechova_fair_2017, barocas_fairness_2023].
+Observational clinical data (e.g., EHR, medical images) inherently contains various biases and researchers must carefully consider sources of bias in any application [genevieve_structural_2020, mccoy_ensuring_2020].
 Additionally, choices made during model training and evaluation can further amplify biases in the data or incorporate new bias, including selecting an inappropriate label (label bias), model (algorithmic bias), and/or evaluation metric or data set (evaluation bias).
@@
 Lastly, during deployment, bias in the data and/or model can be reinforced or introduced when users selectively disregard (dismissal bias) or overly trust (automation bias) a model’s outputs, or if models are inaccessible to certain subpopulation (privilege bias).
-For instance, ML-assisted diagnostic imaging requires that hospitals have the necessary equipment, yet fewer than 20% of rural emergency departments in Canada have in-house computed tomography (CT) scanners [48].
-A review of biases in non-health applications can be found in [43] and [34].
+For instance, ML-assisted diagnostic imaging requires that hospitals have the necessary equipment, yet fewer than 20% of rural emergency departments in Canada have in-house computed tomography (CT) scanners to utilize these advances [davidson_landscape_2024].
+A review of biases outside of health-focused applications can be found in [suresh2021framework] and [mehrabi_survey_2021].
 Dr.
-Leo Celi, a health AI expert, emphasizes that “data bias is the waterloo of health AI” and argues that bias must be a primary consideration throughout the ML pipeline - from problem formulation to model deployment [49].
-Building on existing AI reporting guidelines, [25] introduced a bias evaluation checklist to enable practitioners to systematically and holistically address bias in clinical predictive models [50, 51].
-Bias mitigation strategies and common pitfalls in applying ML in health are further discussed in [52] and [53].
+Leo Celi, a health AI expert, emphasizes that “data bias is the waterloo of health AI.” Bias must be a primary consideration throughout the ML pipeline - from problem formulation to model deployment [tedx_talks_data_2023].
+Building on existing AI reporting guidelines, [wang_bias_2022] introduced a bias evaluation checklist to enable practitioners to systematically and holistically address bias in clinical predictive models [wolff2019probast, liu2020reporting].
+Bias mitigation strategies and common pitfalls in applying ML in health are further discussed in [gichoya2023ai] and [panch_artificial_2019].
 The fairness criteria we introduce in the subsequent sections play an important role in this process.
-The criteria are typically integrated directly into model training or used as evaluative metrics to identify disparities after the model has been trained [54, 55, 56, 57, 41, 58].
-While this review focuses on how to define fairness, strategies for mitigating bias are an equally active and important area of research and have been reviewed from a methodological perspective in [59].
+The criteria are typically integrated directly into model training or used as evaluative metrics to identify disparities after the model has been trained [hardt_equality_2016, menon_cost_2018, pessach_review_2022, brahmbhatt2023towards, mbakwe2023fairness, goethals2024beyond].
+While this review focuses on how to define fairness, strategies for mitigating bias are an equally important area of research and have been reviewed in [hort2023bias].
 (a) Bias in the data
 Type of Bias Definition Example
-Societal/historical Bias The data reflects long-standing societal disparities encoded within the data over time.
-Clinical word embeddings trained on large corpora of text, such as clinical notes from healthcare systems, reflect biases about ethnic minorities [60].
+Societal (or historical) Bias The data reflects long-standing societal disparities encoded within the data over time.
+Clinical word embeddings trained on large corpora of text, such as clinical notes from healthcare systems, reflect biases about ethnic minorities [zhang2020hurtful].
 Selection Bias The data is not representative of the population of interest.
 Data in volunteer-based biobanks does not reflect the general population as participation rates are higher among women, older age groups, and individuals with higher socioeconomic status.
-[61].
+[fry_comparison_2017].
 Measurement Bias The data contains variables that are collected or measured inaccurately.
-Gender identity is often recorded in EHRs without provider education and therefore may not accurately represent an individual’s identity [62].
+Gender identity is often recorded in EHRs without provider education and therefore may not accurately represent an individual’s identity [kronk_transgender_2022].
 Temporal Bias The data captures a specific time period that may not reflect current or future conditions.
-Pediatric mental health visits were lower than expected for individuals with lower socioeconomic status during the first year of the COVID-19 pandemic [63].
-Minority Bias The data lacks adequate representation from the minority group for the model to accurately learn. Most of the data for genetic studies is from European ancestry populations [64].
+Pediatric mental health visits were lower than expected for individuals with lower socioeconomic status during the first year of the COVID-19 pandemic [toulany_sociodemographic_2023].
+Minority Bias The data lacks adequate representation from the minority group for the model to accurately learn.
+Most of the data for genetic studies is from European ancestry populations [martin_clinical_2019].
 Missing Data Bias The data has variables that are incompletely measured.
-Patients from low-income backgrounds have higher rates of missing medical measurements, which is often due to inequities in access to healthcare [65].
+Patients from low-income backgrounds have higher rates of missing medical measurements, which is often due to inequities in access to healthcare [Chapman_Exploring_2022].
 (b) Bias in the model
@@
 Healthcare cost was used as a proxy for healthcare need in a commercial algorithm used to identify patients for high-risk care management programs.
-The algorithm significantly under-identified Black patients for care services as less money is spent on Black patients relative to similarly healthy white patients [16].
+The algorithm significantly under-identified Black patients for care services as less money is spent on Black patients relative to similarly healthy white patients [obermeyer_dissecting_2019].
 Algorithmic Bias Properties of a model and/or its training algorithm create or amplify bias in the data.
-Differentially private prediction models for mortality performed poorly for minority groups [66].
+Differentially private prediction models for mortality performed poorly for minority groups [suriyakumar2021chasing].
 Evaluation Bias An inappropriate choice of benchmark data or metrics are used for evaluation.
-Existing public skin disease AI benchmarks do not have images of biopsy-proven malignancy, the gold standard for disease annotation, on dark skin [67, 68].
+Existing public skin disease AI benchmarks do not have images of biopsy-proven malignancy, the gold standard for disease annotation, on dark skin [daneshjou_lack_2021, daneshjou_disparities_2022].
 (c) Bias in deployment
@@
 A computer-aided diagnosis system that scans mammograms and marks suspicious areas of potential cancer features had lower sensitivity for women aged 40–49 compared to older age groups.
-A clinician using this system can disproportionately miss cancers for this age group [69].
+A clinician using this system can disproportionately miss cancers for this age group [lehman_diagnostic_2015].
 Dismissal Bias Model users ignore model recommendations, often due to frequent false alerts.
-If the same computer-aided diagnosis system also has a lower positive predictive value for women aged 40–49, then clinicians may disregard its recommendations, as they are more likely to be false positives [70].
-Privilege Bias Models are not available or cannot be deployed in all settings, such as parts of the world where medical technology or sensors are unavailable.
-Less than 20% of rural emergency departments in Canada have in-house access to computed tomography (CT) scanners [48].
+If the same computer-aided diagnosis system also has a lower positive predictive value for women aged 40–49, then clinicians may disregard its recommendations, as they are more likely to be false positives [ancker2017effects].
+Privilege Bias Models are not available or cannot be deployed in all settings, such as parts of the world where medical technology is unavailable.
+Less than 20% of rural emergency departments in Canada have in-house access to computed tomography (CT) scanners required for ML-assisted diagnostic imaging [davidson_landscape_2024].
 Table 1: Common biases that arise in machine learning (ML) applications in health and an example from the literature.
@@
 Group fairness criteria are commonly used in health and deem a model as fair if its predictions are similarly accurate or calibrated across a predefined set of groups.
-These groups are most often defined by a protected attribute(s) such as age or race [^2] Sensitive attribute and protected attribute are used interchangeably [71].
-Fairness is also evaluated across groups defined by social determinants of health (e.g., income and education) [72].
-[34].
+These groups are most often defined by a protected attribute(s) such as age or race [^2] Sensitive attribute and protected attribute are used interchangeably [caton_fairness_2024].
+Fairness is also evaluated across groups defined by social determinants of health (e.g., income and education) [mccradden2020ethical].
+[mehrabi_survey_2021].
 Common protected attributes used in the health-focused literature are summarized in Table 2.
-In contrast, individual fairness is a less commonly used framework and requires that the model provide similar predictions to similar individuals based on a user-defined similarity metric [24, 31].
-Lastly, causal fairness criteria utilize causal estimands to quantify unfairness and link observed disparities in model performance to their underlying cause [73, 74].
-Causality-based fairness definitions are particularly attractive for health-focused applications as they enable practitioners to interrogate biases [75].
+In contrast, individual fairness is a less commonly used framework and requires that the model provide similar predictions to similar individuals based on a user-defined similarity metric [dwork_fairness_2012, anderson2024algorithmic].
+Lastly, causal fairness criteria utilize causal estimands to quantify unfairness and link observed disparities in model performance to their underlying cause [plecko2022causal, makhlouf_when_2024].
+Causality-based fairness definitions are particularly attractive for health-focused applications as they enable practitioners to interrogate biases [zhang_causal_2024].
 Attribute Examples
-Age [9, 76, 77, 78]
-Disability [76]
-Marital Status [79]
-National Origin [78]
-Race [9, 79, 76, 77, 80, 78, 33, 25]
-Sex [9, 79, 76, 80, 78]
-Socioeconomic Status [9, 79, 77, 78, 33, 25]
+Age [seyyed-kalantari_chexclusion_2020, garriga_machine_2022, schuch_fairness_2023, yfantidou_accuracy_2023]
+Disability [garriga_machine_2022]
+Marital Status [wang_comparison_2022]
+National Origin [yfantidou_accuracy_2023]
+Race [seyyed-kalantari_chexclusion_2020, wang_comparison_2022, garriga_machine_2022, schuch_fairness_2023, yuan_algorithmic_2023, yfantidou_accuracy_2023, rajkomar_ensuring_2018, wang_bias_2022]
+Sex [seyyed-kalantari_chexclusion_2020, wang_comparison_2022, garriga_machine_2022, yuan_algorithmic_2023, yfantidou_accuracy_2023]
+Socioeconomic Status [seyyed-kalantari_chexclusion_2020, wang_comparison_2022, schuch_fairness_2023, yfantidou_accuracy_2023, rajkomar_ensuring_2018, wang_bias_2022]
 Table 2: Commonly used protected attributes in machine learning for health. Examples of studies that evaluate fairness with respect to each protected attribute.
@@
 Group fairness criteria are oblivious in the sense that they can be entirely inferred from the distribution of the data.
-Individual and causal fairness criteria are non-oblivious criteria as they require specification of a similarity metric and a causal model, respectively.
+Individual and causal fairness criteria are non-oblivious criteria as they require specification of similarity metrics and a causal model, respectively.
 Many fairness criteria cannot be simultaneously satisfied and incompatibilities between different notions of fairness are depicted with a red X.
@@
 Group fairness
-Group fairness criteria require ML models to perform similarly across groups of interest and are the most popular fairness framework in health-focused applications [32].
-The criteria primarily fall into three categories: independence, separation, and sufficiency [81, 17].
+Group fairness criteria require ML models to perform similarly across groups of interest and are the most popular fairness framework in health-focused applications [awasthi2020beyond].
+The criteria primarily fall into three categories: independence, separation, and sufficiency [berk_fairness_2018, barocas_fairness_2023].
 Table 3 provides mathematical definitions of various criteria and their interpretation in the context of an example.
-Our discussion focuses on classification as it is most widely used and well-studied [40, 34, 82].
-Approaches for continuous outcomes, continuous protected attributes, categorical protected attributes that define more than two groups, and multiple protected attributes (i.e., subgroup fairness) are discussed in Supplementary Sections 2.
+Our discussion focuses on classification as it is most widely used and well-studied [mitchell_algorithmic_2021, mehrabi_survey_2021, castelnovo_clarification_2022].
+Approaches for continuous outcomes, continuous protected attributes, categorical protected attributes that define more than two groups, and multiple protected attributes (i.e., subgroup fairness) are discussed in Supplementary Section 2.
 Independence
 Under independence, an ML model is said to be fair if its decisions do not depend on the protected attribute (i.e., $D\perp\!\!\!\perp A$).
-Statistical (or demographic) parity is a common measure of independence that requires that the model classify individuals into the positive class at the same rate in each group [24, 83].
+Statistical (or demographic) parity is a common measure of independence that requires that the model classify individuals into the positive class at the same rate in each group [dwork_fairness_2012, corbett-davies_algorithmic_2017].
 Conditional statistical parity relaxes this concept by requiring the rate of positive classifications to be the same within more granular groups defined by the protected attribute and other relevant factors.
 Independence-based metrics are infrequently used in health-focused applications as the prevalence of clinical outcomes often differs across groups (e.g., heart disease in males and females).
-Enforcing independence may prevent a model from learning a genuine association between the protected attribute and the outcome, which can decrease the overall performance of the model for the entire population [84, 85].
+Enforcing independence may also prevent a model from learning a genuine association between the protected attribute and the outcome, which can decrease the overall performance of the model for the entire population [pfohl2019creating, lei2024inductive].
 Separation
@@
 Common separation-based metrics therefore aim to equalize error rates across the groups, including the false negative rate (FNR, known as equal opportunity), false positive rate (FPR, known as predictive equality), or both (known as equalized odds).
-These metrics have been widely used in health-focused applications as incorrect decisions can have severe consequences.
-The choice of a separation-based metric, however, depends on the specific clinical context.
+Separation-based metrics have been widely used in health-focused applications, although the specific choice of metric depends on the clinical context.
 When false negatives have the most severe consequences, equal opportunity may be preferred.
-For instance, equal opportunity was used in a study of state-of-the-art computer vision models used to detect common diseases from chest X-rays, where a false negative corresponded to incorrectly identifying a patient as not having “no finding” on their X-ray [11].
+For instance, equal opportunity was used in a study of state-of-the-art computer vision models used to detect common diseases from chest X-rays, where a false negative corresponded to incorrectly identifying a patient as not having “no finding” on their X-ray [seyyed-kalantari_underdiagnosis_2021].
 The models had higher FNRs in several intersectional under-served subpopulations, such as Hispanic female patients, potentially resulting in delayed access to care.
-There are also situations where balancing both FNR and FPR is more appropriate and equalized odds should be prioritized [54].
-For instance, in the context of rapid screening tools for COVID-19 used in emergency departments, a high FNR can lead to insufficient monitoring of COVID-19 while a high FPR can result in unnecessary additional testing [86].
-Additional separation-based metrics are detailed in Table 3, including balance for the positive class and balance for the negative class [87].
+There are also situations where balancing both the FNR and FPR is more appropriate and equalized odds should be prioritized [hardt_equality_2016].
+For instance, in the context of rapid screening tools for COVID-19 used in emergency departments, a high FNR can lead to insufficient monitoring of COVID-19 while a high FPR can result in unnecessary additional testing in certain groups [yang2023algorithmic].
+Additional separation-based metrics are detailed in Table 3, including balance for the positive class and balance for the negative class [kleinberg2017inherent].
 Sufficiency
-Sufficiency aims to equalize error rates among individuals with similar decisions [82].
+Sufficiency aims to equalize error rates among individuals with similar decisions [castelnovo_clarification_2022].
 Formally, sufficiency requires that the label does not depend on the protected attribute given the model’s decision ( i.e., $Y\perp\!\!\!\perp A\mid D)$.
-The decision is therefore “sufficient” for predicting the outcome in the sense that it subsumes the protected attributes [17].
-Common sufficiency-based metrics focus on equalizing the positive predictive value (PPV, known as predictive parity), both the PPV and negative predictive value (NPV, known as conditional use accuracy equality), and calibration (known as well-calibration )[45, 81, 87, 88].
-Sufficiency-based metrics are commonly used to evaluate models for disease screening, management, and triage.
-Conditional use accuracy equality is a particularly common metric, as disparities in PPV and NPV can have distinct impacts on patient care [89, 86].
+The decision is therefore “sufficient” for predicting the outcome in the sense that it subsumes the protected attributes [barocas_fairness_2023].
+Common sufficiency-based metrics focus on equalizing the positive predictive value (PPV, known as predictive parity), both the PPV and negative predictive value (NPV, known as conditional use accuracy equality), and calibration (known as well-calibration )[chouldechova_fair_2017, berk_fairness_2018, kleinberg2017inherent, guo_calibration_2017].
+Sufficiency-based metrics are often used to evaluate models for disease screening, management, and triage.
+Conditional use accuracy equality is a particularly common metric, as disparities in PPV and NPV can have distinct impacts on patient care [brandao-de-resende_machine_2023, yang2023algorithmic].
 For example, in COVID-19 triage algorithms, conditional use accuracy seeks to ensure that ICU allocations based on the model’s decisions are determined by medical needs rather than a protected attribute.
@@
 Calibration-based metrics are also frequently applied.
-Well-calibration was used in our introductory example of the commercial risk prediction algorithm used to enroll patients in a high-risk care management program.
+Well-calibration was evaluated in our introductory example of the commercial risk prediction algorithm used to enroll patients in a high-risk care management program.
 In this example, healthcare costs were used as a proxy for health needs in training the model.
-As a result, at equivalent model scores, Black patients were in significantly poorer health than white patients as historically less money had been spent on their healthcare, meaning that they had to be sicker to qualify for the program [16].
+As a result, at equivalent model scores, Black patients were in significantly poorer health than white patients as historically less money had been spent on their healthcare, meaning that they had to be sicker to qualify for the program [obermeyer_dissecting_2019].
 Metric Definition Interpretation Relax-
@@
 Independence-based criteria
-Statistical Parity $P(D=1\mid A=a_{0})=P(D=1\mid A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ In predicting hospitalization and emergency department (ED) visits in heart failure, the predicted probability of a hospitalization or an ED visit is the same for males and females [90].
-Conditional Statistical Parity $P(D=1\mid{\bf Z},A=a_{0})=P(D=1\mid{\bf Z},A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ In predicting hospitalization and emergency department (ED) visits in heart failure, the predicted probability of a hospitalization or an ED visit is the same for males and females after adjusting for age and pre-existing conditions [90].
+Statistical Parity $P(D=1\mid A=a_{0})=P(D=1\mid A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ In predicting hospitalization and emergency department (ED) visits in heart failure, the predicted probability of a hospitalization or an ED visit is the same for males and females [davoudi_fairness_2024].
+Conditional Statistical Parity $P(D=1\mid{\bf Z},A=a_{0})=P(D=1\mid{\bf Z},A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ In predicting hospitalization and ED visits in heart failure, the predicted probability of a hospitalization or an ED visit is the same for males and females after adjusting for age and pre-existing conditions [davoudi_fairness_2024].
 Separation-based criteria
@@
 $P(D=0\mid Y=1,A=a_{0})=P(D=0\mid Y=1,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$
-In predicting asthma exacerbation in children, the rates of false positives (children without exacerbation incorrectly identified as having exacerbation) and false negatives (children with exacerbation incorrectly identified as not having exacerbation) are the same for children from low and high socio-economic classes [91].
-Predictive Equality $P(D=1\mid Y=0,A=a_{0})=P(D=1\mid Y=0,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ A model predicting cardiovascular disease (CVD) diagnoses CVD in males and females without CVD at the same rate [92].
-Equal Opportunity $P(D=0\mid Y=1,A=a_{0})=P(D=0\mid Y=1,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ A model predicting a suspicious finding from a chest X-rays misses findings in males and females at the same rate [93].
-Balance for Positive Class $\mathbb{E}(S\mid Y=1,A=a_{0})=\mathbb{E}(S\mid Y=1,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ In a model used to screen for lung cancer from chest X-rays, the average predicted score of people with lung cancer is equal for males and females [93].
-Balance for Negative Class $\mathbb{E}(S\mid Y=0,A=a_{0})=\mathbb{E}(S\mid Y=0,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ In a model used to screen for lung cancer from chest X-rays, the average scores among males and females without lung cancer are equal [93].
+In predicting asthma exacerbation in children, the rates of false positives (children without exacerbation incorrectly identified as having exacerbation) and false negatives (children with exacerbation incorrectly identified as not having exacerbation) are the same for children from low and high socio-economic classes [juhn2022assessing].
+Predictive Equality $P(D=1\mid Y=0,A=a_{0})=P(D=1\mid Y=0,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ A model predicting cardiovascular disease (CVD) diagnoses CVD in males and females without CVD at the same rate [Vargae002395].
+Equal Opportunity $P(D=0\mid Y=1,A=a_{0})=P(D=0\mid Y=1,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ A model predicting a suspicious finding from a chest X-rays misses findings in males and females at the same rate [ueda_artificial_2021].
+Balance for Positive Class $\mathbb{E}(S\mid Y=1,A=a_{0})=\mathbb{E}(S\mid Y=1,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ In a model used to screen for lung cancer from chest X-rays, the average score of people with lung cancer is equal for males and females [ueda_artificial_2021].
+Balance for Negative Class $\mathbb{E}(S\mid Y=0,A=a_{0})=\mathbb{E}(S\mid Y=0,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ In a model used to screen for lung cancer from chest X-rays, the average score among males and females without lung cancer are equal [ueda_artificial_2021].
 Sufficiency-based criteria
@@
 $P(Y=0\mid D=0,A=a_{0})=P(Y=0\mid D=0,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$
-In predicting hospital readmission, the probability of being readmitted given the model makes that prediction and the probability of not being readmitted given the model makes that prediction is the same for Black and white patients [94].
-Predictive Parity $P(Y=1\mid D=1,A=a_{0})=P(Y=1\mid D=1,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ In predicting hospital readmission, the probability of being readmitted given the model makes that prediction is the same for Black and white patients [94].
-Well Calibration $P(Y=1\mid S=s,A=a)=s\quad\forall a\in\mathcal{A},s\in[0,1]$ In a model used to predict in-hospital mortality, the predicted event rates match the observed event rates at all values of the score for males and females [95].
-Test Fairness $P(Y=1\mid S=s,A=a_{0})=P(Y=1\mid S=s,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A},s\in[0,1]$ In a model used to predict in-hospital mortality, the predicted event rates are the same for males and females at all values of the score [95].
+In predicting hospital readmission, the probability of being readmitted given the model makes that decision and the probability of not being readmitted given the model makes that decision is the same for Black and white patients [raza2023auditing].
+Predictive Parity $P(Y=1\mid D=1,A=a_{0})=P(Y=1\mid D=1,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A}$ In predicting hospital readmission, the probability of being readmitted given the model makes that decision is the same for Black and white patients [raza2023auditing].
+Well Calibration $P(Y=1\mid S=s,A=a)=s\quad\forall a\in\mathcal{A},s\in[0,1]$ In a model used to predict in-hospital mortality, the predicted event rates match the observed event rates at all values of the score for males and females [roosli_2022_peeking].
+Test Fairness $P(Y=1\mid S=s,A=a_{0})=P(Y=1\mid S=s,A=a_{1})\quad\forall a_{0},a_{1}\in\mathcal{A},s\in[0,1]$ In a model used to predict in-hospital mortality, the predicted event rates are the same for males and females at all values of the score [roosli_2022_peeking].
 Other criteria
-Brier Score Parity $\mathbb{E}[(Y-S)^{2}\mid A=a_{0}]=\mathbb{E}[(Y-S)^{2}\mid A=a_{1}],\quad\forall a_{0},a_{1}\in\mathcal{A}$ In a model used to enroll patients into high-risk care management programs, the mean squared error between the score and the label is the same for Black and White patients [16].
-Overall Accuracy Equality $P(D=Y\mid A=a_{0})=P(D=Y\mid A=a_{1})\quad\forall a_{0},a_{1}\in{\cal A}$ In a model used to screen for lung cancer from chest X-rays, the probability of a correct classification (i.e., correctly classifying a patient as having or not having lung cancer) is the same for males and females [93].
-Treatment Equality #$\frac{\#\{D=0,Y=1,A=a_{0}\}}{\#\{D=1,Y=0,A=a_{0}\}}=\frac{\#\{D=0,Y=1,A=a_{1}\}}{\#\{D=1,Y=0,A=a_{1}\}}\quad\forall a_{0},a_{1}\in{\cal A}$ In a model used to screen for lung cancer from chest X-rays, the ratio of the false negatives and false positives is the same for both males and females [93].
+Brier Score Parity $\mathbb{E}[(Y-S)^{2}\mid A=a_{0}]=\mathbb{E}[(Y-S)^{2}\mid A=a_{1}],\quad\forall a_{0},a_{1}\in\mathcal{A}$ In a model used to enroll patients into high-risk care management programs, the mean squared error between the score and the label is the same for Black and White patients [obermeyer_dissecting_2019].
+Overall Accuracy Equality $P(D=Y\mid A=a_{0})=P(D=Y\mid A=a_{1})\quad\forall a_{0},a_{1}\in{\cal A}$ In a model used to screen for lung cancer from chest X-rays, the probability of a correct classification (correctly classifying a patient as having or not having lung cancer) is the same for males and females [ueda_artificial_2021].
+Treatment Equality #$\frac{\#\{D=0,Y=1,A=a_{0}\}}{\#\{D=1,Y=0,A=a_{0}\}}=\frac{\#\{D=0,Y=1,A=a_{1}\}}{\#\{D=1,Y=0,A=a_{1}\}}\quad\forall a_{0},a_{1}\in{\cal A}$ In a model used to screen for lung cancer from chest X-rays, the ratio of the false negatives and false positives is the same for both males and females [ueda_artificial_2021].
 Table 3: Common group fairness criteria.
 Mathematical definitions of group fairness criteria, their interpretation in the context of an example from the literature, and an indication of whether the criterion is a relaxation of independence ($D\perp\!\!\!\perp A$), separation ($D\perp\!\!\!\perp A\mid Y$), or sufficiency ($Y\perp\!\!\!\perp A\mid D$).
-Symbols: #=\#=# = number of, $P=$ probability, $E=$ expected value.
+Symbols: #=#absent\#=# = number of, $P=$ probability, $E=$ expected value.
 Notations: $Y:$ outcome, ${\bf X}:$ features used for model training, $A:$ protected attribute that takes value in the set $\mathcal{A}$, $S:$ model score, $D:$ model classification based on thresholding $S$, ${\bf Z}$: additional set of features.
@@
 Independence, sufficiency, and separation provide different perspectives on what it means for a model to be fair.
-Except under highly restrictive conditions, it is not possible for an algorithm to fulfill all criteria simultaneously [54, 45, 96].
-It is therefore critical for healthcare researchers to choose which group fairness considerations are most relevant to the specific context.
+Except under highly restrictive conditions, it is not possible for an algorithm to fulfill all criteria simultaneously [hardt_equality_2016, chouldechova_fair_2017, corbett-davies_measure_2023].
+It is therefore critical for healthcare researchers to choose which group fairness considerations are most relevant to their context.
 More specifically, the following pairs of criteria are incompatible in the sense that they cannot generally be simultaneously satisfied: independence and sufficiency, independence and separation, and separation and sufficiency.
@@
 In classification problems, this means the probability of being in the positive class is the same across groups - a condition often violated in clinical contexts, where disease prevalence may vary by factors like age or sex.
-Supplementary Section 3 provides additional mathematical details related to the incompatibilities and also introduces approximate fairness, which allows for small deviations in group fairness metrics to ease these incompatibilities.
-In Supplementary Section 4, we further explore the practical implications of these incompatibilities through a case study using an openly available EHR dataset.
+Supplementary Section 3 provides additional mathematical details related to the incompatibilities and also introduces approximate fairness, which allows for small deviations in group fairness metrics to relax these incompatibilities.
+In Supplementary Section 4, we also explore the practical implications of these incompatibilities through a case study using an openly available EHR dataset.
 Individual fairness
-In contrast to group fairness that targets the average performance of a model across groups, individual fairness ensures that “like cases are treated alike”, an idea grounded in Aristotle’s consistency principles [24, 97].
+In contrast to group fairness that targets the average performance of a model across groups, individual fairness ensures that “like cases are treated alike”, an idea grounded in Aristotle’s consistency principles [dwork_fairness_2012, fleisher_whats_2021].
 One of the first works on individual fairness, fairness through awareness (FTA), formalized the notion of similarity by defining metrics that measure the distance between individuals and their predictions.
 Fairness is achieved if the distance between the prediction for any two individuals is at most the distance between them.
-A special case of FTA is when the model yields identical decisions for two individuals with all but their protected attributes being the same, a property known as causal discrimination [98].
-The implementation of FTA-based approaches critically depends on the choice of distance metrics defined on both predicted outcomes and individuals.
-For predictions, the choice varies by the type of outcome under study.
-For example, the Euclidean distance is used for logit transformed predicted probabilities in classification while the absolute difference of the hazard function has been used for survival analysis [99, 100].
-However, the metric used to measure the similarity between individuals has been a focus of ongoing research as it encodes exactly how individuals are similar, which relies on an “awareness” of the context [^4] For consistency index proposed by [101], the distance metric for predictions are absolute difference, but it still relies on the choice of metric to define $k$ nearest neighborhood..
+A special case of FTA is when the model yields identical decisions for two individuals with all but their protected attributes being the same, a property known as causal discrimination [galhotra_fairness_2017].
+The implementation of FTA-based approaches critically depends on the metrics used to define similarity between individuals and their predictions.
+For predictions, the choice of metric depends on the type of outcome.
+For example, the Euclidean distance has been used for logit transformed predicted probabilities in classification while the absolute difference of the hazard function has been used for survival analysis [keya2021equitable, rahman2022fair].
+However, the metric used to measure the similarity between individuals has been a focus of ongoing research as it determines how individuals are comparable, which relies on an “awareness” of the context [^4] For consistency index proposed by [zemel_learning_2013], the distance metric for predictions are absolute difference, but it still relies on the choice of metric to define $k$ nearest neighborhood..
 In the original FTA paper, the authors suggest leaving the choice of metric to domain experts.
-As there is currently no agreed upon standard, particularly within health settings, various strategies have been proposed to learn the metric from available data [102, 103, 104].
-A scoping review on individual fairness in healthcare provides a comprehensive list of individual fairness methods and corresponding software [31].
-The review highlights that individual fairness is just emerging within health-focused applications, likely due to the relative nascency of individual fairness methods developed within the ML literature and the difficult in appropriately quantifying similarity among individuals.
-It is important to note that FTA is a counter approach to fairness through unawareness (FTU), which is intended to be a catch all solution to prevent bias by not explicitly including protected attributes into modeling [105] [^5] Note that fairness through awareness implies causal discrimination, but not vice versa..
+As there is currently no agreed upon standard, particularly within health settings, various strategies have been proposed to learn the metric from available data [kim2018fairness, ilvento2019metric, mukherjee2020two].
+A scoping review on individual fairness in healthcare provides a comprehensive list of individual fairness methods and corresponding software [anderson2024algorithmic].
+The review highlights that individual fairness is just emerging within health-focused applications, likely due to the relative nascency of individual fairness methods developed within the ML literature and the difficult in appropriately quantifying similarity across individuals.
+It is important to note that FTA is a counter approach to fairness through unawareness (FTU), which is intended to be a catch all solution to prevent bias by not explicitly including protected attributes into modeling [kusner_counterfactual_2017] [^5] Note that fairness through awareness implies causal discrimination, but not vice versa..
 However, removing protected attributes from a model warrants careful consideration in health-focused applications.
-First, protected attributes may serve as critical predictors for the outcome of interest, such as age in sepsis or cardiovascular disease prediction and race in cancer screening models [106, 107, 108, 109].
+First, protected attributes may serve as critical predictors for the outcome of interest, such as age in sepsis or cardiovascular disease prediction and race in cancer screening models [damen2016prediction, kim_sex_2018, chicco2020survival, cirillo2020sex].
 Excluding these predictors can diminish the overall predictive accuracy and impact all individuals adversely or even lead to bias against the majority group.
@@
 Simply removing protected attributes from the model does not prevent the model from inferring them from other attributes.
-This is evident in a recent study that found ML models can infer race from medical images [110].
+This is evident in a recent study that found ML models can infer race from medical images [gichoya_ai_2022].
 Finally, the quality of the other predictors included in the model may vary across the protected groups (e.g., rates of missing data, measurement error).
-For instance, in cancer risk prediction, family history serves as an important predictor for disease risk, but was shown to be less reliably documented for Black participants in EHR and therefore less predictive for models developed with these data.
-[111] found that including race as a predictor significantly improved model performance compared to race-blind algorithms for colorectal cancer.
-FTU is therefore not always ethical or achievable [112].
+For instance, in cancer risk prediction, family history serves as an important predictor for disease risk, but was shown to be less reliably documented for Black participants in self-reported family history data and therefore less predictive for models developed with these data [kupfer_racial_2006, chavez-yenter_association_2022, zink_race_2024].
+[zink_race_2024] found that including race as a predictor significantly improved model performance compared to race-blind algorithms for colorectal cancer.
+FTU is therefore not always ethical, achievable, or desirable [banja2023fairness].
 Lastly, individual fairness is motivated by the inherent weakness in group fairness criteria that only consider average model performance within groups.
 That is, there are situations in which group fairness can be satisfied, but individuals within a group can be discriminated against.
-The incompatibilities among individual and group fairness have been examined from a theoretical perspective in [113, 114, 82, 115].
-For example, in predicting ICU admission, if the rate of admission differs significantly across protected attributes, the predictions must be adjusted to achieve statistical parity and obtain an equal rate of positive predictions across the groups.
+The incompatibilities among individual and group fairness have been examined from a theoretical perspective in [friedler_impossibility_2016, binns_apparent_2020, castelnovo_clarification_2022, xu_compatibility_2024].
+For example, in predicting ICU admission, if the rate of admission differs significantly across protected attributes, the predictions must be adjusted to achieve statistical parity and obtain an equal rate of positive classifications across the groups.
 This can cause individuals with similar clinical profiles and medical needs to receive different predictions based solely on their protected attributes, which contradicts the principle of individual fairness.
-However, individual fairness implies statistical parity when the Wasserstein distance between two groups’ protected attribute distributions is small [24, 116].
+However, individual fairness implies statistical parity when the Wasserstein distance between two groups’ protected attribute distributions is small [dwork_fairness_2012, zhou_group_2022].
 Causal fairness
-We conclude with a brief overview of causal fairness, an increasingly prominent topic in ML that warrants a dedicated review within the health context [117, 118, 119, 120, 121, 31, 75].
-As the name suggests, causal fairness focuses on understanding the causal relationships between protected attributes and a model’s decisions [30].
+We conclude with a brief overview of causal fairness, an increasingly prominent topic in ML that warrants a dedicated review within the health context [loftus2018causal, fawkes2022selection, makhlouf2022survey, carey2022causal, nilforoshan_causal_2022, anderson2024algorithmic, zhang_causal_2024].
+As the name suggests, causal fairness focuses on understanding the causal relationships between protected attributes and a model’s decisions [plecko_causal_2022].
 Although causality-based fairness approaches are especially valuable in health applications for mitigating confounding effects and disentangling mechanisms of bias, the use of causal fairness is health is in its early stages.
-We introduce one way of incorporating causality into fairness through counterfactuals, which are simply “what-if” statements [122, 123].
-In fairness context, the counterfactual is often the unobserved model’s decision that would have happened if the protected attribute had been different [105, 118, 124].
-One of the first counterfactual-based fairness criteria, simply called (individual) counterfactual fairness, deems a model fair if the observed and counterfactual predictions align for an individual when their protected attributed is changed, given all other variables of interest are held constant [105] [^6] Counterfactual fairness is also viewed as an individual fairness criteria as it is grounded in achieving similar treatment for similar individuals[105]..
+We introduce one way of incorporating causality into fairness through counterfactuals, which are simply “what-if” statements .
+In fairness context, the counterfactual is often the unobserved model’s decision that would have happened if the protected attribute had been different [kusner_counterfactual_2017, fawkes2022selection, bynum2024new].
+One of the first counterfactual-based fairness criteria, simply called (individual) counterfactual fairness, deems a model fair if the observed and counterfactual predictions align for an individual when their protected attributed is changed, given all other variables of interest are held constant [kusner_counterfactual_2017] [^6] Counterfactual fairness is also viewed as an individual fairness criteria as it is grounded in achieving similar treatment for similar individuals[kusner_counterfactual_2017]..
 To better explain this concept, consider an over-simplified view of the world in predicting 30-day hospital readmission.
-The directed acyclic graph (DAG) in Figure 3 illustrates that the model’s decision for 30-day hospital readmission is directly related to previous diagnoses, age, and prescriptions, while race indirectly influences the decision through its impact on access to healthcare, which in turn influences historical diagnoses [125].
+The directed acyclic graph (DAG) in Figure 3 illustrates that the model’s decision for 30-day hospital readmission is directly related to previous diagnoses, age, and prescriptions, while race indirectly influences the decision through its impact on access to healthcare, which in turn influences historical diagnoses [pfohl_counterfactual_2019].
 When predicting 30-day hospital readmission, counterfactual fairness is satisfied if a Black patient receives the same prediction they would have received if they were counterfactually of a different race, given all other variables are held constant.
-The ceteris paribus condition here is crucial, as in reality healthcare utilization may differ across racial communities, which could influence the rate of 30-day hospital readmission in practice.
-A challenge to causal fairness criteria is that they require a working causal model, the assumptions of which cannot be verified from observational data [105, 126, 118].
-Other quantities from the causal inference literature have since been proposed, such as counterfactual parity which measures a population-level, rather than a individual-level, causal effect of the protected attribute on the model’s decision [127, 40, 128].
-Conditional counterfactual fairness is a weaker criteria that considers the total effect of the protected attribute on the decision conditional on a set of variables of interest [105].
+The ceteris paribus condition here is crucial, as in reality healthcare utilization may differ within different communities defined by race, which could influence the rate of 30-day hospital readmission in practice.
+A challenge to causal fairness criteria is that they require a working causal model, the assumptions of which cannot be verified from observational data [kusner_counterfactual_2017, salimi2019interventional, fawkes2022selection].
+Numerous other estimands from the causal inference literature have been proposed, such as counterfactual parity which measures a population-level, rather than a individual-level, causal effect of the protected attribute on the model’s decision [pearl2009causality, mitchell_algorithmic_2021, alves_survey_2023].
+Conditional counterfactual fairness is a weaker criteria that considers the total effect of the protected attribute on the decision conditional on a set of variables of interest [kusner_counterfactual_2017].
 In our previous example, this means that for each group of individuals with identical characteristics (e.g., those with the same historical diagnosis, age and prescriptions), the 30-day hospital readmission does not change if their race were counterfactually altered.
-Another criteria used in health-focused applications is principle fairness, which is a population-level fairness criteria that evaluates decisions among individuals who would be similarly impacted by a model’s decision [129].
-Formally, a model satisfies principle fairness if the model’s decision is independent of the protected attribute in principle stratum, which are subgroups with the same values of the counterfactual decision under all possible decisions [129].
-More concretely, in our example, principle fairness is achieved if Black and white people have an equal chance of being flagged for 30-day readmission if they have the same likelihood of 30-day readmission (i.e., they belong to the same principal strata).
-These definitions, along with extensions of the previously introduced group fairness measures — such as counterfactual equalized odds, counterfactual predictive parity, and conditional principal fairness — are formally introduced in Supplementary Section 6.
-Further work on causal fairness, such as path-specific fairness, direct and indirect discrimination, and interventional fairness, are detailed in [130, 131, 132, 133, 134, 135, 136, 126, 137].
+Another criteria used in health-focused applications is principle fairness, which is a population-level fairness criteria that evaluates decisions among individuals who would be similarly impacted the decisions [imai_principal_2023].
+Formally, a model satisfies principle fairness if the model’s decision is independent of the protected attribute in principle stratum, which are subgroups with the same values of the counterfactual decision under all possible decisions [imai_principal_2023].
+More concretely, in our example, principle fairness is achieved if Black and white people have an equal chance of being flagged for 30-day readmission if they have a similar likelihood of 30-day readmission (i.e., they belong to the same principal strata).
+These definitions, along with extensions of the previously introduced group fairness measures — such as counterfactual equalized odds, counterfactual predictive parity, and conditional principal fairness — are mathematically defined in Supplementary Section 6.
+Further work on causal fairness, such as path-specific fairness, direct and indirect discrimination, and interventional fairness, are detailed in [chiappa_path-specific_2019, nabi_fair_2018, wu_pc-fairness_2019, zhang_causal_2017, zhang_mitigating_2018, ogburn_comment_2019, wang_blessings_2019, salimi2019interventional, zuo_interventional_2024].
 Figure 3: Directed acyclic graph (DAG) for predicting 30-day hospital readmission.
@@
 This review explores why models are unfair and various ways fairness has been defined in ML for health.
-Despite substantial progress in recognizing the importance of bias and in applying group, individual, and causal-based criteria in health applications, there is no consensus on how to appropriately quantify fairness [138, 139].
+Despite substantial progress in recognizing the importance of bias and in applying group, individual, and causal-based criteria in health applications, there is no consensus on how to appropriately quantify fairness [zong_medfair_2023, gichoya_equity_2021].
 Each framework comes with unique challenges.
@@
 However, specifying reliable causal models can be challenging in health contexts.
-Balancing these different frameworks involves weighing the benefits of group-level equity, individual needs, and causal insights [140].
+Balancing these different frameworks involves weighing the benefits of group-level equity, individual needs, and causal insights [schrouff_algorithmic_2022].
 Researchers must remain informed about the full range of available methods and work within interdisciplinary teams (e.g., data scientists, clinicians, ethicists) to make appropriate decisions for their specific application.
-It is critical to note that a quantitative understanding of fairness is just one component of operationalizing fairness in ML for health.
-Establishing common principles, guidelines, and standards that emphasize fairness across the entire model development process is essential, as simply assessing fairness after model training is insufficient [141].
-Moreover, most existing fairness definitions were developed outside the health context and do not account for underlying inequities or potential harms to underserved populations [47].
-There is an urgent need for context-aware fairness criteria tailored to the unique challenges of health applications [142].
-Additionally, health data present inherent challenges, making it difficult - or sometimes impossible - to measure fairness accurately [143, 144, 145, 146, 147, 148].
-Various protected attributes (e.g., attributes such as religion, gender identity, and sexual orientation that are missing from Table 2) are often unavailable or inaccurately measured in health datasets due to socio-cultural challenges and lack of standardized data collection practices [149, 62, 150].
-While by no means a remedy for inadequate data, there has been recent progress in fairness approaches that operate with limited data on protected attributes that may prove useful in health applications [151, 152].
-To date, however, many protected attributes have been omitted from fairness considerations, which underscores the necessity of future methodological work to acknowledge “a multiplicity of considerations, from privacy preservation, context sensitivity and process fairness, to an awareness of sociotechnical impact and the increasingly important role of inclusive and participatory research processes” [153].
+A quantitative understanding of fairness is just one component of operationalizing fairness in ML for health.
+Establishing common principles, guidelines, and standards that emphasize fairness throughout model development is essential, as simply assessing fairness after model training is insufficient [mccradden_whats_2023].
+Health data present inherent and unique challenges, making it difficult - or sometimes impossible - to measure fairness accurately [wick2019unlocking, ji2020can, choi2020fair, zhang2021fairness].
+Most existing fairness definitions were developed outside the health context and do not account for underlying inequities or potential harms to underserved populations [mccoy_ensuring_2020].
+There is a need for context-aware fairness criteria tailored for health applications [lalor_should_2024].
+Additionally, various protected attributes (e.g., attributes such as religion, gender identity, and sexual orientation that are missing from Table 2) are often unavailable or inaccurately measured in health datasets due to sociocultural issues and lack of standardized data collection practices [kronk_transgender_2022, ruberg2020data].
+While by no means a remedy for inadequate data, there has been recent progress in fairness approaches that operate with limited data on protected attributes that may prove useful in health applications [zhang_fairlisa_2023, roy_fairness_2024].
+To date, however, many protected attributes have been omitted from fairness considerations, which underscores the necessity of future methodological work to acknowledge “a multiplicity of considerations, from privacy preservation, context sensitivity and process fairness, to an awareness of sociotechnical impact and the increasingly important role of inclusive and participatory research processes” [tomasev2021fairness].
 We anticipate future work in these aforementioned directions, particularly in strategies focused on data collection and promoting data equity in health contexts, to establish a strong foundation for fair decision-making.
@@
 All authors provided valuable feedback on the manuscript.
+[…]